# Extended symmetry analysis of (1+2)-dimensional fine Kolmogorov backward equation

### v1

@@
[…]
${}^{\dagger}$epartment of Mathematics and Statistics, Memorial University of Newfoundland,
@@
[…]
${}^{\ddagger}$ Institute of Mathematics of NAS of Ukraine, 3 Tereshchenkivska Str., 01024 Kyiv, Ukraine
@@ 1 Introduction
The equations $\mathcal{F}_{0}$ and $\mathcal{F}_{2}$ of the form (1) with $\beta=0$ and $\beta=2$ are distinguished by their symmetry properties not only in the narrow class $\mathcal{F}$ but also in the very wide class $\bar{\mathcal{F}}$. More specifically, the essential Lie invariance algebra of the equation $\mathcal{F}_{0}$ is eight-dimensional, which is the maximum of such dimensions among equations from the class $\bar{\mathcal{F}}$. Moreover, possessing the essential Lie invariance algebra of dimension eight is an attribute of the equation $\mathcal{F}_{0}$ that singles it out, up to the point equivalence, within the class $\bar{\mathcal{F}}$. This is why the second equation in the class $\mathcal{F}$ with eight-dimensional essential Lie invariance algebra, $\mathcal{F}_{5}$, is necessarily equivalent to $\mathcal{F}_{0}$ with respect to point transformations. The essential Lie invariance algebra of the equation $\mathcal{F}_{2}$ is five-dimensional and nonsolvable, and modulo the point equivalence, it is the only equation from the class $\bar{\mathcal{F}}$ with this property. Within the class $\mathcal{F}$, there are only two equations with five-dimensional essential Lie invariance algebras, $\mathcal{F}_{2}$ and $\mathcal{F}_{3}$, and they are similar with respect to a point transformation. Because of this, we refer to the equations (1) with $\beta=0$ and $\beta=2$ as the remarkable Fokker–Planck equation and the fine Kolmogorov backward equation, respectively. The equation $\mathcal{F}_{0}$ is the canonical form of the ultraparabolic Fokker–Planck equations with constant diffusion and drift coefficients, whose study was initiated by Kolmogorov in [11] as mentioned above. Moreover, it is shown in [2] that the Kolmogorov equations describing the price of the geometric and arithmetic average strike call option are reduced by point transformations to Eqs. (3.4) and (4.5) [2], which coincide with $\mathcal{F}_{0}$ and $\mathcal{F}_{2}$ up to alternating the sign of $y$, respectively.
Due to its special form, the remarkable Fokker–Planck equation can also be considered as a Kolmogorov backward equation. Its extended symmetry analysis was carried out in [13], where we computed its point-symmetry pseudogroup using the direct method and analyzed its structure, which allowed us to list one- and two-dimensional subalgebras of its essential and maximal Lie invariance algebra with a view towards the classification of the Lie reductions and finding Lie invariant solutions. Moreover, we have carried out peculiar generalized reductions that allowed us to construct three wide families of the solutions of the remarkable Fokker–Planck equation parameterized by an arbitrary finite number of solutions of the linear (1+1)-dimensional heat equation.
@@ 2 Lie invariance algebra
The maximal Lie invariance $\mathfrak{g}$ algebra of (2) is spanned by the vector fields
$\mathcal{P}^{y}=\partial_{y},\quad\mathcal{D}=x\partial_{x}+y\partial_{y},\quad\mathcal{K}=2xy\partial_{x}+y^{2}\partial_{y}-ux\partial_{u},$
@@ 3 Point symmetry pseudogroup
$\begin{split}&\tilde{t}=t+\lambda,\quad\tilde{x}=\frac{\alpha\delta-\beta\gamma}{(\gamma y+\delta)^{2}}x,\quad\tilde{y}=\frac{\alpha y+\beta}{\gamma y+\delta},\\[4.30554pt] &\tilde{u}=\sigma\exp\left(-\frac{\gamma}{\gamma y+\delta}x\right)\big{(}u+f(t,x,y)\big{)}\end{split}$ (5)
@@
We integrate the system (3), where both source and target arbitrary-element tuples are equal to $\bar{\theta}_{2}$, with respect to the parameter-functions $T$, $X$, $Y$ and $U$. The first two equations (4c) take the form
@@
This decomposition correspond to the Lie algebra decomposition $\mathfrak{g}^{\rm ess}=\mathfrak{f}\oplus\mathfrak{r}\simeq{\rm sl}(2,\mathbb{R})\oplus 2A_{1}$.
@@ 4 Classification of subalgebras
Subalgebras of a Lie algebra $\mathfrak{g}$ can be classified up to various equivalence relations. However, from the algebraic perspective, the most interesting and useful is the classification modulo the action of the inner automorphism group ${\rm Inn}(\mathfrak{g})$ of $\mathfrak{g}$. Moreover, when $\mathfrak{g}$ is a real (resp. complex) finite-dimensional Lie algebra, this action coincides with the adjoint action of the corresponding real (resp. complex) simply-connected Lie group on $\mathfrak{g}$. Another natural equivalence for the subalgebra classification is generated by the action of the entire automorphism group ${\rm Aut}(\mathfrak{g})$ of $\mathfrak{g}$. Note that it often happens that there is no big difference between the above two equivalences or, in more precise algebraic terms, ${\rm Inn}(\mathfrak{g})={\rm Aut}(\mathfrak{g})$ or at least the index of ${\rm Inn}(\mathfrak{g})$ as a subgroup of ${\rm Aut}(\mathfrak{g})$ is finite and small.
@@
$\hphantom{\rm 2D\colon}\quad\smash{\mathfrak{s}_{2.6}^{\mu,\mu^{\prime}}}=\langle\mathcal{P}^{y}+\mathcal{K}+\mu I,\mathcal{P}^{t}+\mu^{\prime}\mathcal{I}\rangle,\quad\mathfrak{s}_{2.7}^{\mu}=\langle\mathcal{P}^{y}+\mathcal{K}+\mu\mathcal{P}^{t},\mathcal{I}\rangle,$
@@ 5 Lie reductions
Throughout this section, the subscripts 1 and 2 of the dependent variable $w$ denote the derivatives with respect to $z_{1}$ and $z_{2}$, respectively. We use the complex label $d.i^{*}$ to mark the Lie reduction corresponding to the subalgebra $\mathfrak{s}_{d.i}^{*}$ from Corollary 7, where $d$, $i$ and *** are the dimension of this subalgebra, its number in the list of $d$-dimensional subalgebras of the algebra $\mathfrak{g}^{\rm ess}$, and the values of subalgebra parameters for a subalgebra family, respectively. We omit the superscript in the label whenever it is not essential.
@@
Table 1: $G$-inequivalent codimension-one Lie reductions of the equation (2)
@@
[…]
Here $\delta^{\prime}\in\{0,1\}$, $\mu\in\mathbb{R}$, $\nu\geqslant 0$, $\nu^{\prime}>0$, and $\mu\geqslant 0$ in Cases 1.5 and 1.7 with $\nu=0$; $w=w(z_{1},z_{2})$ is the new unknown function of the new independent variables $(z_{1},z_{2})$.
@@
$\tilde{\mathfrak{a}}_{1.1}^{\mu}=\langle\partial_{z_{1}},z_{1}\partial_{z_{1}}+z_{2}\partial_{z_{2}},2z_{1}z_{2}\partial_{z_{2}}+z_{1}^{2}\partial_{z_{1}}-z_{2}w\partial_{w},w\partial_{w}\rangle,\quad\tilde{\mathfrak{a}}_{1.4}^{0}=\langle\partial_{z_{1}},z_{2}\partial_{z_{2}},w\partial_{w}\rangle$
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=\ln|z_{2}|,\ \ \tilde{w}={\rm e}^{z_{1}/4}|z_{2}|^{-1/2}w:\ \ \tilde{w}_{1}=\tilde{w}_{22}-\delta\tilde{\varepsilon}{\rm e}^{\tilde{z}_{2}}\tilde{w}$ with $\tilde{\varepsilon}=\mathop{\rm sgn}\nolimits z_{2}$;
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=2\sqrt{1-\nu^{\prime}z_{2}}+\ln\left|\dfrac{\sqrt{1-\nu^{\prime}z_{2}}-1}{\sqrt{1-\nu^{\prime}z_{2}}+1}\right|,\ \ \tilde{w}=\dfrac{{\rm e}^{z_{2}/2}}{|z_{2}|^{1/2}}(1-\nu^{\prime}z_{2})^{1/4}w:$
@@
$\tilde{z}_{1}=-z_{1},\ \ \tilde{z}_{2}=2\sqrt{\nu^{\prime}z_{2}-1}-2\arctan\sqrt{\nu^{\prime}z_{2}-1},\ \ \tilde{w}=\dfrac{{\rm e}^{z_{2}/2}}{z_{2}^{1/2}}(\nu^{\prime}z_{2}-1)^{1/4}w:$
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=\ln|z_{2}|,\ \ \tilde{w}={\rm e}^{z_{1}/4+z_{2}/2}|z_{2}|^{-1/2}w:\\ \tilde{w}_{1}=\tilde{w}_{22}-(\tilde{\varepsilon}\nu{\rm e}^{\tilde{z}_{2}}+\frac{1}{4}{\rm e}^{2\tilde{z}_{2}})\tilde{w}$  with  $\tilde{\varepsilon}=\mathop{\rm sgn}\nolimits z_{2}$;
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=2\sqrt{1-\nu^{\prime}z_{2}}+\ln\left|\dfrac{\sqrt{1-\nu^{\prime}z_{2}}-1}{\sqrt{1-\nu^{\prime}z_{2}}+1}\right|,\ \ \tilde{w}=\dfrac{(1-\nu^{\prime}z_{2})^{1/4}}{|z_{2}|^{1/2}}w$
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=2\sqrt{\nu^{\prime}z_{2}-1}-2\arctan\sqrt{\nu^{\prime}z_{2}-1},\ \ \tilde{w}=\dfrac{(\nu^{\prime}z_{2}-1)^{1/4}}{z_{2}^{1/2}}w$
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=\ln|z_{2}|,\ \ \tilde{w}=\dfrac{{\rm e}^{z_{1}/4}}{|z_{2}|^{1/2}}w:\ \ \tilde{w}_{1}=\tilde{w}_{22}+({\rm e}^{2\tilde{z}_{2}}-\tilde{\varepsilon}\nu{\rm e}^{\tilde{z}_{2}})\tilde{w}$  with  $\tilde{\varepsilon}:=\mathop{\rm sgn}\nolimits z_{2}$.
@@
$\bullet\quad u=|x|^{1/4}\theta(\varepsilon y,2\sqrt{|x|}),$ (6)
@@ 6 Generating new solutions from known ones
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@
Serhii D. Koval†‡ and Roman O. Popovych§‡
Department of Mathematics and Statistics, Memorial University of Newfoundland,
@@
§ Mathematical Institute, Silesian University in Opava, Na Rybníčku 1, 746 01 Opava, Czech Republic
 Institute of Mathematics of NAS of Ukraine, 3 Tereshchenkivska Str., 01024 Kyiv, Ukraine
@@ 1 Introduction
The equations $\mathcal{F}_{0}$ and $\mathcal{F}_{2}$ of the form (1) with $\beta=0$ and $\beta=2$ are distinguished by their symmetry properties not only in the narrow class $\mathcal{F}$ but also in the very wide class $\bar{\mathcal{F}}$. More specifically, the essential Lie invariance algebra of the equation $\mathcal{F}_{0}$ is eight-dimensional, which is the maximum of such dimensions among equations from the class $\bar{\mathcal{F}}$. Moreover, possessing the essential Lie invariance algebra of dimension eight is an attribute of the equation $\mathcal{F}_{0}$ that singles it out, up to the point equivalence, within the class $\bar{\mathcal{F}}$. This is why the second equation in the class $\mathcal{F}$ with eight-dimensional essential Lie invariance algebra, $\mathcal{F}_{5}$, is necessarily equivalent to $\mathcal{F}_{0}$ with respect to point transformations. The essential Lie invariance algebra of the equation $\mathcal{F}_{2}$ is five-dimensional and nonsolvable, and modulo the point equivalence, it is the only equation from the class $\bar{\mathcal{F}}$ with this property. Within the class $\mathcal{F}$, there are only two equations with five-dimensional essential Lie invariance algebras, $\mathcal{F}_{2}$ and $\mathcal{F}_{3}$, and they are similar with respect to a point transformation. Because of this, we refer to the equations (1) with $\beta=0$ and $\beta=2$ as the remarkable Fokker–Planck equation and the fine Kolmogorov backward equation, respectively. The equation $\mathcal{F}_{0}$ is the canonical form of the ultraparabolic Fokker–Planck equations with constant diffusion and drift coefficients $f$ and $g$, whose study was initiated by Kolmogorov in [11] as mentioned above. Moreover, it is shown in [2] that the Kolmogorov equations describing the price of the geometric and arithmetic average strike call option are reduced by point transformations to Eqs. (3.4) and (4.5) [2], which coincide with $\mathcal{F}_{0}$ and $\mathcal{F}_{2}$ up to alternating the sign of $y$, respectively.
Due to its special form, the remarkable Fokker–Planck equation can also be considered as a Kolmogorov backward equation. Its extended symmetry analysis was carried out in [13], where we computed its point-symmetry pseudogroup using the direct method and analyzed its structure, which allowed us to list one- and two-dimensional subalgebras of its essential and maximal Lie invariance algebra with a view toward the classification of the Lie reductions and finding Lie invariant solutions. Moreover, we have carried out peculiar generalized reductions that allowed us to construct three wide families of the solutions of the remarkable Fokker–Planck equation parameterized by an arbitrary finite number of solutions of the linear (1+1)-dimensional heat equation.
@@ 2 Lie invariance algebra
The maximal Lie invariance algebra $\mathfrak{g}$ of (2) is spanned by the vector fields
$\mathcal{P}^{y}=\partial_{y},\quad\mathcal{D}=x\partial_{x}+y\partial_{y},\quad\mathcal{K}=2xy\partial_{x}+y^{2}\partial_{y}-xu\partial_{u},$
@@ 3 Point symmetry pseudogroup
$\begin{split}&\tilde{t}=t+\lambda,\quad\tilde{x}=\frac{\alpha\delta-\beta\gamma}{(\gamma y+\delta)^{2}}x,\quad\tilde{y}=\frac{\alpha y+\beta}{\gamma y+\delta},\\[4.30554pt] &\tilde{u}=\sigma\exp\left(-\frac{\gamma}{\gamma y+\delta}x\right)\big{(}u+f(t,x,y)\big{)},\end{split}$ (5)
@@
We integrate the system (4), where both source and target arbitrary-element tuples are equal to $\bar{\theta}_{2}$, with respect to the parameter-functions $T$, $X$, $Y$ and $U$. The first two equations from (4c) take the form
@@
This decomposition corresponds to the Lie algebra decomposition $\mathfrak{g}^{\rm ess}=\mathfrak{f}\oplus\mathfrak{r}\simeq{\rm sl}(2,\mathbb{R})\oplus 2A_{1}$.
@@ 4 Classification of subalgebras
Subalgebras of a Lie algebra $\mathfrak{g}$ can be classified up to various equivalence relations. However, from the algebraic perspective, the most interesting and useful is the classification modulo the action of the inner automorphism group ${\rm Inn}(\mathfrak{g})$ of $\mathfrak{g}$. Moreover, when $\mathfrak{g}$ is a real (resp. complex) finite-dimensional Lie algebra, this action coincides with the adjoint action of the corresponding real (resp. complex) simply connected Lie group on $\mathfrak{g}$. Another natural equivalence for the subalgebra classification is generated by the action of the entire automorphism group ${\rm Aut}(\mathfrak{g})$ of $\mathfrak{g}$. Note that it often happens that there is no big difference between the above two equivalences or, in more precise algebraic terms, ${\rm Inn}(\mathfrak{g})={\rm Aut}(\mathfrak{g})$ or at least the index of ${\rm Inn}(\mathfrak{g})$ as a subgroup of ${\rm Aut}(\mathfrak{g})$ is finite and small.
@@
$\hphantom{\rm 2D\colon}\quad\smash{\mathfrak{s}_{2.6}^{\mu,\mu^{\prime}}}=\langle\mathcal{P}^{y}+\mathcal{K}+\mu\mathcal{I},\mathcal{P}^{t}+\mu^{\prime}\mathcal{I}\rangle,\quad\mathfrak{s}_{2.7}^{\mu}=\langle\mathcal{P}^{y}+\mathcal{K}+\mu\mathcal{P}^{t},\mathcal{I}\rangle,$
@@ 5 Lie reductions
Throughout this section, the subscripts 1 and 2 of the dependent variable $w$ denote the derivatives with respect to $z_{1}$ and $z_{2}$, respectively. We use the complex label $d.i^{*}$ to mark the Lie reduction corresponding to the subalgebra $\mathfrak{s}_{d.i}^{*}$ from Corollary 7, where $d$, $i$ and * are the dimension of this subalgebra, its number in the list of $d$-dimensional subalgebras of the algebra $\mathfrak{g}^{\rm ess}$, and the values of subalgebra parameters for a subalgebra family, respectively. We omit the superscript in the label whenever it is not essential.
@@
Table 1: $G^{\rm ess}$-inequivalent codimension-one Lie reductions of the equation (2)
@@
1.5ν,μ $|y|^{\mu}w$ $t-\nu\ln|y|$ $x/y$ $(1-\nu z_{2})w_{1}=z_{2}^{2}w_{22}+z_{2}^{2}w_{2}-\mu z_{2}w$
1.7ν,μ ${\rm e}^{\mu\arctan(y)-\frac{xy}{y^{2}+1}}w$ $t-\nu\arctan y$ $\dfrac{x}{y^{2}+1}$ $(1-\nu z_{2})w_{1}=z_{2}^{2}w_{22}+z_{2}(z_{2}-\mu)w$
Here $\delta^{\prime}\in\{0,1\}$, $\mu\in\mathbb{R}$, $\nu\geqslant 0$, and $\mu\geqslant 0$ in Cases 1.5 and 1.7 with $\nu=0$; $w=w(z_{1},z_{2})$ is the new unknown function of the new independent variables $(z_{1},z_{2})$.
@@
$\tilde{\mathfrak{a}}_{1.1}^{\mu}=\langle\partial_{z_{1}},z_{1}\partial_{z_{1}}+z_{2}\partial_{z_{2}},2z_{1}z_{2}\partial_{z_{2}}+z_{1}^{2}\partial_{z_{1}}-z_{2}w\partial_{w},w\partial_{w}\rangle,\quad\tilde{\mathfrak{a}}_{1.4}^{0}=\langle\partial_{z_{1}},z_{2}\partial_{z_{2}},w\partial_{w}\rangle,$
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=\ln|z_{2}|,\ \ \tilde{w}={\rm e}^{z_{1}/4}|z_{2}|^{-1/2}w,\ \ \tilde{w}_{1}=\tilde{w}_{22}-\delta\tilde{\varepsilon}{\rm e}^{\tilde{z}_{2}}\tilde{w}$ with $\tilde{\varepsilon}=\mathop{\rm sgn}\nolimits z_{2}$;
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=2\sqrt{1-\nu^{\prime}z_{2}}+\ln\left|\dfrac{\sqrt{1-\nu^{\prime}z_{2}}-1}{\sqrt{1-\nu^{\prime}z_{2}}+1}\right|,\ \ \tilde{w}=\dfrac{{\rm e}^{z_{2}/2}}{|z_{2}|^{1/2}}(1-\nu^{\prime}z_{2})^{1/4}w$,
@@
$\tilde{z}_{1}=-z_{1},\ \ \tilde{z}_{2}=2\sqrt{\nu^{\prime}z_{2}-1}-2\arctan\sqrt{\nu^{\prime}z_{2}-1},\ \ \tilde{w}=\dfrac{{\rm e}^{z_{2}/2}}{z_{2}^{1/2}}(\nu^{\prime}z_{2}-1)^{1/4}w$,
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=\ln|z_{2}|,\ \ \tilde{w}={\rm e}^{z_{1}/4+z_{2}/2}|z_{2}|^{-1/2}w,\\ \tilde{w}_{1}=\tilde{w}_{22}-(\tilde{\varepsilon}\nu{\rm e}^{\tilde{z}_{2}}+\frac{1}{4}{\rm e}^{2\tilde{z}_{2}})\tilde{w}$  with  $\tilde{\varepsilon}=\mathop{\rm sgn}\nolimits z_{2}$;
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=2\sqrt{1-\nu^{\prime}z_{2}}+\ln\left|\dfrac{\sqrt{1-\nu^{\prime}z_{2}}-1}{\sqrt{1-\nu^{\prime}z_{2}}+1}\right|,\ \ \tilde{w}=\dfrac{(1-\nu^{\prime}z_{2})^{1/4}}{|z_{2}|^{1/2}}w$,
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=2\sqrt{\nu^{\prime}z_{2}-1}-2\arctan\sqrt{\nu^{\prime}z_{2}-1},\ \ \tilde{w}=\dfrac{(\nu^{\prime}z_{2}-1)^{1/4}}{z_{2}^{1/2}}w$,
@@
$\tilde{z}_{1}=z_{1},\ \ \tilde{z}_{2}=\ln|z_{2}|,\ \ \tilde{w}=\dfrac{{\rm e}^{z_{1}/4}}{|z_{2}|^{1/2}}w,\ \ \tilde{w}_{1}=\tilde{w}_{22}+({\rm e}^{2\tilde{z}_{2}}-\tilde{\varepsilon}\nu{\rm e}^{\tilde{z}_{2}})\tilde{w}$  with  $\tilde{\varepsilon}:=\mathop{\rm sgn}\nolimits z_{2}$.
@@
$\bullet\quad u={\rm e}^{-\frac{3}{16}t}|x|^{1/4}\theta(\varepsilon y,2\sqrt{|x|}),$ (6)
@@ 6 Generating new solutions from known ones
1.1μ. For the above reason, the Lie-symmetry operator $\hat{\mathrm{P}}^{t}-\mu$, which corresponds to the Lie-symmetry vector field $-\mathcal{P}^{t}-\mu\mathcal{I}$ of (2), commutes with any monomial $\mathbf{Q}^{\alpha}$. Hence the action by any element from $\Upsilon_{\mathfrak{f}}$ maps the set of $\mathfrak{s}_{1.1}^{\mu}$-invariant solutions to itself. This implies that the solution families (6) and (8) of the equation (2) cannot be extended using the generation of its solutions by the action of its Lie-symmetry operators according to (10). Another manifestation of the above phenomenon is that each Lie-symmetry vector field of (6) induces a Lie-symmetry vector field of reduced equation 1.1μ. Hence the algebra $\Upsilon_{\mathfrak{f}}$ induces the entire algebra generated by Lie-symmetry operators of reduced equation 1.1μ if $\mu\neq-\frac{3}{16}$ and a proper subalgebra of the latter algebra otherwise.
@@
1.31,μ. For the same reason as above, the operator $\mathrm{P}^{y}+\hat{\mathrm{P}}^{t}-\mu$ commutes with any polynomial in $\mathrm{P}^{y}$ and $\hat{\mathrm{P}}^{t}$. This is why the action by such polynomials maps the space of $\mathfrak{s}_{1.3}^{1,\mu}$-invariant solutions to itself. It can be shown that the centralizer of $\mathrm{P}^{y}$ in $\Upsilon_{\mathfrak{f}}$ and thus the centralizer of $\mathrm{P}^{y}+\hat{\mathrm{P}}^{t}-\mu$ in $\Upsilon_{\mathfrak{f}}$ coincide with the space of these polynomials.
@@
1.5ν,μ. Recall that this case corresponds to the two subalgebra families listed in Corollary 7, $\{\mathfrak{s}_{1.5}^{\nu^{\prime},\mu}\}$ and $\{\mathfrak{s}_{1.6}^{\nu}\}$, see the beginning of Section 5. Since $\mathrm{K}\mathrm{D}=(\mathrm{D}-1)\mathrm{K}$ and $\mathrm{P}^{y}\mathrm{D}=(\mathrm{D}+1)\mathrm{P}^{y}$, we have
@@
1.7ν,μ. We can only state that the action by each polynomial in $\mathrm{P}^{y}+\mathrm{K}$ and $\hat{\mathrm{P}}^{t}$ maps the space of $\mathfrak{s}_{1.7}^{\nu,\mu}$-invariant solutions to itself since such polynomials commute with the operator $\mathrm{P}^{y}+\mathrm{K}+\nu\hat{\mathrm{P}}^{t}-\mu$.